OpenStudy (anonymous):

Find the area of a trapezoid with a height of 6 m and bases of 13 m and 26 m. 45 square meters 58.5 square meters 234 square meters 117 square meters 2028 square meters

OpenStudy (anonymous):

The formula for the area of a trapezoid is: \[A = \frac{1}{2}(b_1+b_2)h\] so just plug in your values and we get: \[A = \frac{1}{2}(26+13)(6) = 117\]

OpenStudy (anonymous):